OpenStudy (anonymous):

The formula for the area of a rectangle is A=lw. What is the length of the rectangle if the area is 48cm^2 and the width is 6cm?

OpenStudy (anonymous):

ANSWER CHOICES ARE

OpenStudy (hba):

l=A/w I=48/6 I=..

OpenStudy (hba):

You can do the rest i guess :)

OpenStudy (anonymous):

A. 6 cm B. 8 cm C. 10 cm D. 12 cm

OpenStudy (anonymous):

is it squared or to the second power

OpenStudy (anonymous):

oh nvm its area

OpenStudy (anonymous):

so yea squared or second power

OpenStudy (anonymous):

got it it is 8 cm^2

OpenStudy (hba):

l=A/w I=48/6 I=.. Just divide

OpenStudy (hba):

No see the units too.

OpenStudy (anonymous):

yea 8 cm squared

OpenStudy (hba):